# Reformulating Sequential Recommendation: Learning Dynamic User Interests with Content-enriched Language Modeling

### v2

@@ 4.2 Results and Analyses
Table 2: Recommendation performance comparison of different recommendation models over three datasets, in which the best results are in bold and the second best results are underlined.
Dataset Metric Model
Caser GRU4Rec SASRec BERT4Rec FDSA MV-RNN SASRec-C LANCER
MovieLens Recall@5 0.0225 0.0737 0.0967 0.0776 0.0901 0.0853 0.0944 0.1086
Recall@10 0.0330 0.1316 0.1589 0.1391 0.1407 0.1399 0.1454 0.1616
NDCG@5 0.0151 0.0450 0.0624 0.0466 0.0595 0.0527 0.0618 0.0712
NDCG@10 0.0185 0.0635 0.0823 0.0664 0.0757 0.0703 0.0781 0.0884
MIND Recall@5 0.0312 0.0981 0.1087 0.0853 0.1063 0.1024 0.1107 0.1168
Recall@10 0.0575 0.1569 0.1707 0.1359 0.1728 0.1619 0.1753 0.1675
NDCG@5 0.0181 0.0628 0.0714 0.0544 0.0682 0.0656 0.0723 0.0766
NDCG@10 0.0265 0.0816 0.0913 0.0706 0.0896 0.0847 0.0930 0.0930
Goodreads Recall@5 0.1058 0.1871 0.2043 0.1892 0.1713 0.1872 0.2045 0.2121
Recall@10 0.1452 0.2472 0.2641 0.2487 0.2351 0.2486 0.2663 0.2673
NDCG@5 0.0770 0.1394 0.1536 0.1391 0.1195 0.1385 0.1533 0.1600
NDCG@10 0.0898 0.1587 0.1729 0.1583 0.1400 0.1583 0.1733 0.1779
In terms of sequential recommendation, our experimental results are shown in Table 2. Based on the results, we can conclude that LANCER achieves better performance than previous models in movie, news and book recommendations. Therefore, it can be seen that LANCER can integrate various textual features of items well, utilizing the powerful text understanding and generation capabilities of PLMs, as well as incorporating knowledge and item features in the recommendation field, bringing more personalized, personified, and precise recommendations for each user.
Additionally, we notice several findings regarding our model. First, LANCER performs more prominently on MovieLens, which we attribute to two factors: 1) the relatively small size of the dataset allows our model to leverage the extensive knowledge contained in the pre-trained model, unleashing its full potential. 2) Movie titles provide limited information, and incorporating more information (such as genre, synopsis, director, etc.) and user movie-watching behavior can lead to more precise reasoning and better recommendations. Second, we notice that the improvement in the @5 metrics is greater than that in the @10 metrics, which we attribute to the generation capabilities of the language model. Generative language models often perform better in the first few generations, but the quality of generations tends to decline as the number of generations required increases [23, 8, 4]. However, we use a generative self-regressive model that can better fit the characteristics of sequential data, which is also why we choose GPT-2 as our backbone.
@@
Figure 3: Case study of three users from the MovieLens-1M dataset in terms of top-5 ranked predictions generated by LANCER. The texts in green indicate predictions that map to the target.
@@
In Figure 3, we select three user examples from the MovieLens-1M dataset and present the top 5 output texts directly generated by the proposed LANCER. It is evident that our model is capable of generating recommendations that align reasonably well with user preferences. Furthermore, we observe that model outputs are realistic movie titles in almost all scenarios, even when the model fails to accurately recommend the next item. Case 2 is among the few examples where the model output do not exactly match a movie title in the database. In this case, the original model output is mapped to Carnosaur 2, which is the correct next item. In Case 3, although the target item does not appear in the top 5 predictions, several predicted movies share overlapping genres with the target. Both Welcome to the Dollhouse and Hero are "Comedy|Drama" movies, and The Godfather also lies in the "Drama" category.
Ablation Study.
Table 3: Ablation results for various LANCER models and P5-single on three different datasets, in which the best results are in bold. R and N denote Recall and NDCG respectively.
Model MovieLens MIND Goodreads
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
P5-single 0.0503 0.0659 0.0370 0.0421 0.0369 0.0588 0.0244 0.0314 0.1358 0.1778 0.1019 0.1154
LANCER-id 0.0715 0.1030 0.0465 0.0567 0.0931 0.1328 0.0620 0.0748 0.1582 0.1734 0.1304 0.1354
LANCER-bare 0.0263 0.0459 0.0130 0.0193 0.0974 0.1511 0.0551 0.0724 0.1492 0.1822 0.0836 0.0944
LANCER-title 0.0998 0.1493 0.0654 0.0814 0.1104 0.1556 0.0730 0.0878 0.1992 0.2267 0.1527 0.1618
LANCER-full 0.1086 0.1616 0.0712 0.0884 0.1168 0.1675 0.0766 0.0930 0.2121 0.2673 0.1600 0.1779
In this section, we mainly discuss the effect of incorporating item content information into sequential recommendation and the impact of knowledge prompts and reasoning prompts. We design four LANCER recommendation methods: LANCER-full represents the same setting as the main experiment, LANCER-title uses only item title text without considering other richer content information, LANCER-id uses only item IDs as text representation, and LANCER-bare removes all templates and prompts and simply concatenates item titles for recommendation. We also implement P5-single, which is trained only on single sequential recommendation tasks due to dataset limitations. We adopt the same 13 different prompts for sequential recommendation as in [6]. We report Recall and NDCG (@5, @10) and the experimental results are shown in Table 3.
According to the experimental results, it can be observed that the LANCER-full incorporating domain knowledge information and item content information achieve the best results on all three datasets. Additionally, we can draw the following conclusions: 1) using IDs for recommendation is weaker than those using textual information, which we believe is due to the language model’s better understanding of well-defined textual information. 2) The recommendation results of P5-single are weaker than LANCER-id, which demonstrates the advantages of autoregressive models in sequential recommendation tasks. 3) LANCER-title performs best on the MIND dataset, followed by Goodreads, and worst on MovieLens. We believe this is due to the different domain characteristics and data types of the three datasets. The MIND dataset, as a news recommendation dataset, already contains sufficient information in the news title to summarize the specific content, and people often determine their interest by browsing news titles. The MovieLens dataset often has short movie titles with limited information, serving only as a sales tool to attract people to watch the movie or seek more information, so people can not determine their interest based solely on a movie title.
Study of Decoding Strategy.
In this section, we explore the impact of different decoding strategies and mapping methods on recommendation performance. We adopt four different decoding strategies: Sample, which generates a token sequence by sampling and then calculates the top $n$ items with the highest cosine similarity in the item database; Beam Sample, which simultaneously generates $n$ token sequences, and for each generated result, calculates the item with the highest cosine similarity to obtain $n$ recommendation results; Greedy Search, which is similar to Sample but adopts a greedy search to generate the token sequence; and Beam Search, which is similar to Beam Sample but uses beam search to generate the token sequence. The experimental results are shown in Figure 4. We can observe that the beam search strategy outperforms the sampling strategy, and mapping $n$ generated results to the top item with the highest cosine similarity is better than mapping a single generated result to the top $n$ items with the highest cosine similarity. We believe that this is because in the high-dimensional vector space composed of all texts, the item text representations are only sparse points, so the generated text can easily match the correct item in the top-1 similarity.
Figure 4: The impact of different decoding strategy on recommendation results.
Study of Input Templates.
In this section, we explore the impact of using different natural language templates to package user and item interaction sequences on the recommendation results. First, we fix the text instruction of the input text and try using different texts to separate each item. The results as shown in Figure 4(a). Afterwards, we fix the separator between each item and modify the text instruction of the input text. We conduct experiments on the MovieLens-1M dataset and compare the results as shown in Figure 4(b).
According to the experimental results, we find that different natural language templates have a certain impact on the recommendation results. Firstly, incorporating text instructions at the beginning of interaction sequences can enhance the performance. Furthermore, language models show a preference
(a) Separators.
(b) Text instructions.
Figure 5: The impact of different natural language templates on sequential recommendation results.
for logical expressions. More logical expressions are favored by the language model, while some inferior templates may even yield negative effects when compared to recommendations without any additional text modification. This observation can be attributed to the characteristics of PLMs, which include human expression and behavioral logic in their training corpus. Therefore, prompts with templates that are more in line with human cognition will perform better.
@@ 5 Conclusion
In conclusion, we propose LANCER, a novel paradigm that utilizes a language model as a sequential recommender, effectively integrating domain-specific knowledge and general knowledge of the PLM to learn dynamic user interests. Our approach employs a knowledge prompt to facilitate the acquisition of item content and a reasoning prompt to enable the integration of both domain-specific knowledge and user behavior for personalized recommendation. We highlight the potential of our approach in bridging the gap between language models and recommender systems and provide experimental results to support its effectiveness in incorporating item content understanding. We hope that this work will inspire future researchers to delve into content-based modeling for recommendation using language models, thereby fostering iterative advancements in the field of recommender systems.

### v4

@@ 3.2 Results and Analyses
In terms of sequential recommendation, our experimental results in Table 2 demonstrate that LANCER outperforms previous models across movie, news, and book recommendations. Therefore, it can be seen that LANCER can integrate various textual features of items well, utilizing the powerful text understanding and generation capabilities of PLMs. Additionally, LANCER incorporates knowledge and item features in the recommendation domain, delivering more personalized and precise recommendations to users.
Additionally, we notice observe several notable points in our results. First, LANCER performs more prominently on MovieLens, which we attribute to two factors: 1) the relatively small size of the dataset allows our model to leverage the extensive knowledge contained in the pre-trained model, unleashing its full potential. 2) Movie titles offer limited information, and further integration of details like genre, synopsis, director, etc., along with user movie-watching behavior, could enhance reasoning and recommendations. Second, we notice that the improvement in the @5 metrics is greater than that in the @10 metrics, which we attribute to the generation capabilities of the language model. Generative LMs often perform better in the first few generations, but the quality of generations tends to decline as the number of generations increases. The reason we choose GPT-2 as our backbone is that generative autoregressive models better fit the characteristics of sequential data.
@@
Figure 3: Case study of three users from the MovieLens dataset in terms of top-5 ranked predictions. The texts in green indicate predictions that map to the target.
@@
In Figure 3, we select three user examples from the MovieLens-1M dataset and present the top 5 output texts directly generated by the proposed LANCER. It is evident that our model is capable of generating recommendations that align reasonably well with user preferences. Furthermore, we observe that model outputs are realistic movie titles in almost all scenarios, even when the model fails to accurately recommend the next item. Case 2 is among the few examples where the model output do not exactly match a movie title in the database. In this case, the original model output is mapped to Carnosaur 2, which is the correct next item. In Case 3, although the target item does not appear in the top 5 predictions, several predicted movies share overlapping genres with the target. Both Welcome to the Dollhouse and Hero are Comedy|Drama movies, and The Godfather also lies in the Drama category.
@@ 4 Conclusion
In conclusion, we propose LANCER, which utilizes language models as sequential recommender systems, effectively integrating domain-specific knowledge and general knowledge of the PLM to learn dynamic user interests. Our approach employs a knowledge prompt to facilitate the acquisition of item content and a reasoning prompt to enable the integration of both domain-specific knowledge and user behavior for personalized recommendation. We highlight the potential of our approach in bridging the gap between language models and recommender systems and provide experimental results to support its effectiveness in incorporating item content understanding.
4.0.1 Acknowledgements.
This research was supported by grants from the National Natural Science Foundation of China (Grants No. 62337001, 623B1020) and the Fundamental Research Funds for the Central Universities, the Fundamental Research Funds for the Central Universities and the Joint Fund for Medical Artificial Intelligence (Grant No. MAI2022C004).